OpenStudy (anonymous):

Choose the correct answer. Anita and Joelle bowl together and their combined total score for one game was 425 points. Anita’s score was 40 less than twice Joelle’s. What were their scores? (Points : 1) Joelle: 270, Anita: 155 Joelle: 180, Anita: 245 Joelle: 140, Anita: 240 Joelle: 155, Anita: 270

OpenStudy (anonymous):

gaga oh lala

OpenStudy (anonymous):

lol @succk.my.success

OpenStudy (texaschic101):

x + y = 425 --> x = 425 - y now sub 425 - y in for x y = 2x - 40 y = 2(425 - y) - 40 -- distribute through the parenthesis y = 850 - 2y - 40 2y + y = 850 - 40 3y = 810 y = 810/3 y = 270 now sub 270 in for y in either of the equations to find x x + y = 425 x + 270 = 425 x = 425 - 270 x = 155 anita's score (y) = 270 joelle's score (x) = 155

OpenStudy (blackops2luvr):

the answer is \(\frak {\color{blue} {u} }\)

OpenStudy (anonymous):

do wat the texas girl said

OpenStudy (anonymous):

wow thank u for walking me trough this i mostly understand the beginning towards the middle i kept gtting messed up at the end

OpenStudy (anonymous):